OpenStudy (anonymous):

if logbase12 27 =a then the value of logbase6 16 is

OpenStudy (imtiaz7):

Remember that you can change a log to any base you want, using this rule: log[base b](a) = log[base c](a) / log[base c](b) So here you have log[base 12](27) = a log[base 6](27) / log[base 6](12) = a Now we need to get "log[base 6](16)" somewhere out of this: log[base 6](27) = a log[base 6](12) log[base 6](27) = a log[base 6](4*3) log[base 6](27) = a ( log[base 6](4) + log[base 6](3) ) log[base 6](27) = a ( log[base 6](16^(1/2)) + log[base 6](3) ) log[base 6](27) = a ( (1/2)log[base 6](16) + log[base 6](3) ) Now just solve this so that "log[base 6](16)" is on one side.

OpenStudy (imtiaz7):

@23ayesha1994 no problem.

OpenStudy (dumbcow):

also 27 = 3^3 you can get answer in terms of "a" and log_base6 (3)

OpenStudy (anonymous):

no imtiaz thankz for help

OpenStudy (imtiaz7):

ok so i need help from u can u help me?

OpenStudy (anonymous):

ya sure

OpenStudy (imtiaz7):

http://openstudy.com/updates/51f61d72e4b00daf471cacaa in this q